OpenStudy (anonymous):

Anyone know "new" like fractions? Fractions" 2/ over 3 and 1/ over 3: Denominator:15 ?

OpenStudy (agreene):

I dont really know what a "new" like fraction is, but I assume your question to be asking for a fraction with denominator of 15 that is equal to these you've given. So, I'll help you with that. So, to get a denominator of 15 from 2/3 we need to multiply top and bottom by some number, which will equal 15 on bottom, 3*5= 15 so, we do that to both top and bottom: 2 * 5 10 ---- = -- 3 * 5 15 2/3 = 10/15 Now you just need to multiply top and bottom of 1/3 to get your second answer (since it has the same denominator).

OpenStudy (anonymous):

thanks! ugh dont like math